during the current year, music company's piano division reports sales of 750,000 and operating assets of 375,000. what is the sales margin that the piano division must earn in order to achieve a return on investment of 15%?
Answer #1

ROI = Margin * turnover

Turnover = Sales / Average operating assets

= 750,000 / 375,000 = 2

15 = Margin * 2

Margin = 15/2 = 7.50%